Day shift, Monday June 7, 1999

Herman B. White (hwhite@fnal.gov)
Mon, 07 Jun 1999 15:44:13 -0500 (CDT)

Shift Summary for day shift, 6/7/99

On Shift: Herman White, Yau Wai Wah

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): ~4-5 cosmics

Accounting of lost time: LCW problems,and power supply problems


========================================================================
Runs ended this shift:
Run 13413 - muons

Tapes brought to FCC this shift: 0
(including tapes for online split)


========================================================================
Accesses during shift:
A number of accesses were made during the repair period.
All accesses were controlled.



========================================================================
Description of Shift:

8:00 - Brian Lavoy and HBW inspect He Bags (particularly #5). Some small
leaks were repaired near the hose inlet of HeBag#5. Found the
flow meters for HeBag #1 and HeBag #2 FULLY opened. After
correcting, this flow rate was abated and set to
nominal values. It remains stable.
Note the pressure on the tub trailer at 1600 psi.


08:45 - Water leak being investigated at NM2. The NM3s power supply
needs to be replaced. Assorted repairs in the accelerator leads
to a estimated 18 hour shutdown.

9:00 - End run 13413 (muon straight through run, ~1 M events)


14:05 - Rick Coleman and HBW make access to increase the 90-degree
monitor 3-fold coincidence unit output width from ~9ns to ~13ns.

14:55 - E. Blucher and S.E. Turner, swap out dpmt 1800, that cured a wide
ped rms

Limited activity this shift. Waiting for beam..